OpenStudy (babynini):

Approximate, to nearest 10', the solutions of the equation in the interval [0 (degrees), 360 (Degrees)) 12 tan^2t-5tant -2 = 0

OpenStudy (babynini):

@optiquest @zepdrix Sorry you guuuys

zepdrix (zepdrix):

ooo crapper... i dont remember how to do `degree minutes seconds`. Are you allowed to use a calculator on this? I would assume so..

OpenStudy (babynini):

Yeah, so i just need the solutions and then i'm good to do the minute thing

zepdrix (zepdrix):

oh ok cool

OpenStudy (babynini):

so factoring tant(12tant-5)-2 = 0

zepdrix (zepdrix):

no no no silly billy.

OpenStudy (babynini):

aw -.-

zepdrix (zepdrix):

you have a quadratic equation. how do we solve quadratics? it either factors nicely, or it doesn't (in which case we use our quadratic function to solve).

OpenStudy (babynini):

so how do I solve it?

zepdrix (zepdrix):

your problem really looks like this:\[\Large\rm 12x^2-5x-2=0\]Where \(\Large\rm x=\tan t\)

zepdrix (zepdrix):

Hmm won't factor nicely ;p quadratic function time.

zepdrix (zepdrix):

plug in all the bits and pieces, tell me what numbers you get

OpenStudy (babynini):

dangit...quad formula was a long time ago lol

zepdrix (zepdrix):

do it -_- nao!

OpenStudy (babynini):

I doing eeet I doing eet

OpenStudy (babynini):

\[(-5\pm \sqrt{5^{2}-4(12)(2)})/2(12)(2)\]

OpenStudy (babynini):

right?

zepdrix (zepdrix):

too many 2's in the bottom and hold up, your c value is -2, not 2.

OpenStudy (babynini):

and it's -5 too right? :|

zepdrix (zepdrix):

you gotta learn to use the equation tool sometime XD lol

zepdrix (zepdrix):

oh yes, -5 as well

OpenStudy (babynini):

soo ( -5 (+/-) 11)/48

zepdrix (zepdrix):

no, still too many 2's in the bottom

zepdrix (zepdrix):

bottom is 2(a), yah?

zepdrix (zepdrix):

i think you accidentally put 2(a)(c)

OpenStudy (babynini):

isn't it 2(a)(c)?

zepdrix (zepdrix):

nooo :O

OpenStudy (babynini):

ooh...

OpenStudy (babynini):

so -5(+/-)11/24

zepdrix (zepdrix):

yay good \c:/

zepdrix (zepdrix):

Simplify it further to 2 values

Nnesha (nnesha):

hmm \[\huge\rm 12 x^2 - 5 x-2 =0\] tan t = x find two number if you multiply them you should get product of ac and if you add or subtract them you should get middle term which is -5 heehe how about headphone method ??;3

zepdrix (zepdrix):

Nooo, they're fractions >.< that sounds like a lot of work. *grumble grumble*

Nnesha (nnesha):

<----

OpenStudy (babynini):

wait what? the answers are 6/24, and -16/24 right?

OpenStudy (babynini):

oh take note that this time they want the answers within degrees

zepdrix (zepdrix):

oh oh we made a mistake back here: -5(+/-)11/24 should have been 5(+/-)11/24

OpenStudy (babynini):

oopsies

zepdrix (zepdrix):

the first part is -(b) plugging in your b gives you -(-5)

zepdrix (zepdrix):

So i guess you get 16/24 and -6/24 ya? simplify furtherrrrrrr -_-

OpenStudy (babynini):

0.66 and -0.25

zepdrix (zepdrix):

So we have:\[\Large\rm x=\frac{2}{3},\qquad -\frac{1}{4}\]So,\[\Large\rm \tan t=\frac{2}{3},\qquad -\frac{1}{4}\]

zepdrix (zepdrix):

To solve for t, we'll have to use our inverse function.

zepdrix (zepdrix):

\[\Large\rm \tan t=\frac{2}{3}\]So,\[\Large\rm \tan^{-1}\left(\frac{2}{3}\right)=t\]

zepdrix (zepdrix):

make sure you're in degree mode

OpenStudy (babynini):

33.690

zepdrix (zepdrix):

good good good. tangent is period in 180 degrees, so we get the same 2/3 if we add 180 degrees to our angle, ya?

zepdrix (zepdrix):

So our first two solutions for our angle t are, 33.690 and 213.690

zepdrix (zepdrix):

How bout the others?\[\Large\rm \tan^{-1}\left(-\frac{1}{4}\right)=t\]

OpenStudy (babynini):

hrm my calculator doesn't like that it says syntax error -.-

OpenStudy (babynini):

it doesn't like the negative in front of the 1/4 but is it - 14.036?

zepdrix (zepdrix):

are you using the subtraction sign instead of the (-) button by mistake?

OpenStudy (babynini):

....meybee

zepdrix (zepdrix):

yesss -14 degrees. that doesn't lie in our interval though, right? so we'll add 180 degrees to find the next one.

OpenStudy (babynini):

165.963

OpenStudy (babynini):

and 345.963

zepdrix (zepdrix):

mmm ok good

zepdrix (zepdrix):

i THINK we did these correctly. i'm trying to see if i can check our work on a graph or something >.< sec

OpenStudy (babynini):

kks, ah, you always go the extra mile. :D

zepdrix (zepdrix):

ehhh im not really sure on this one >.< we'll just go with this, it's probably right. you understand how to convert your angles to degrees, minutes, seconds from that point?

OpenStudy (babynini):

oh shooo i forgot to do that haha THANK YOU

zepdrix (zepdrix):

yay team \c:/

OpenStudy (babynini):

yayaa